OpenStudy (anonymous):

how do i do the problem (2+5) squared?

OpenStudy (jhannybean):

is it \(\sqrt {2+5}\)?

OpenStudy (anonymous):

no, it is (2+5) and then the 2 outside the parentheses to sqaure it. i know the answer is 49 im just confused as to how i get that i cant figure it out

OpenStudy (anonymous):

\[(a+b)^{2} = a^2 + 2ab + b^{2}\] useful to know when you have unknown variables, like (x +2)^2. the equation holds for any a and b

OpenStudy (jhannybean):

Hmm....\(\sqrt{(2+5)^2}\)? that would be \[\large \sqrt{7^2}=\sqrt{49}\]

OpenStudy (jhannybean):

and i'll let you figure that out :)

OpenStudy (anonymous):

ah thank you!

OpenStudy (jhannybean):

no problem, and good explaining Euler

OpenStudy (anonymous):

4 squared plus 8 squared plus 2 squared is 32, but how do get that?

OpenStudy (anonymous):

i got 80 adding all of those together so i am confused

OpenStudy (anonymous):

PEMDAS

OpenStudy (jhannybean):

\[\large 4^2+8^2+2^2\] This does not = 32. (4)(4)+(8)(8)=(2)(2)

OpenStudy (jhannybean):

multiply all those out and then add them all together.

OpenStudy (anonymous):

the back of my book says its 32 so the book is totally wrong

OpenStudy (jhannybean):

Maybe :\ unless there is more to the problem?

OpenStudy (anonymous):

no there is not any more to the problem

OpenStudy (jhannybean):

Yeah then i believe the book is wrong. That cannot possibly be 32.

OpenStudy (jhannybean):

Anything else? :)

OpenStudy (anonymous):

i know lol my book is trying to confuse me

OpenStudy (jhannybean):

xD